OpenStudy (anonymous):

5(2x - 7) = 33 - 3(6x + 4) . Be sure to show all work. You should have a minimum of 7 lines of work.

OpenStudy (amistre64):

lol .... riiggghhhttt

OpenStudy (amistre64):

if you were to show ALL the work, youd have at a minimum 7 pages

OpenStudy (angela210793):

10x-35=33-18x-12 10x+18x=21+35 28x=56 x=56/28 x=8/4 x=2

OpenStudy (angela210793):

oops i have only 6....but we can do this 10x-35=33-18x-12 10x+18x=21+35 28x=56 x=56/28 x=(7*8)/(7*4) x=8/4 x=2

OpenStudy (ybarrap):

5 (2 x-7) = 33-3 (6 x+4) Expand out terms on both sides: 10 x-35 = 21-18 x Add (18 x+35) to both sides: 28 x = 56 Divide both sides by 28: x = 2

OpenStudy (anonymous):

thank u so much guyz!

OpenStudy (ybarrap):

np

OpenStudy (amistre64):

5 (2x - 7) = 33 - 3 (6x + 4) ; x = 2

OpenStudy (angela210793):



OpenStudy (anonymous):